OpenStudy (waheguru):

guys i relly need help on this question FAST To compare the efficiency of a microwave with the efficiency of an electric kettle, Sara placed 500 mL of water into each. Sara’s teacher helped her calculate the energy required to boil the water, which was 168 kJ. The kettle, which had a power rating of 1.5 kW, took 132 s to boil the water. The microwave had a power rating of 1.2 kW and took 280 s to boil the water. Which is more efficient?

OpenStudy (waheguru):

i know efficancy is found energy out divided energy in x 100 but i cant find the output energy

OpenStudy (jamesj):

Power = (change in energy)/(change in time) = Energy/Time. Hence Energy = Power x Time So the kettle expended 1.5 kW x 132 s = 198 kJ and the microwave 1.2 kW x 280 s = 336 kJ See the answer now?

OpenStudy (waheguru):

so is the kettle more efficant??

OpenStudy (waheguru):

and how come the 168kj part is not needed

OpenStudy (jamesj):

Yes, because it expended less energy than the microwave, both of which ended up doing the same effective work.

OpenStudy (jamesj):

You can calculate the efficiency explicitly with the 168kJ if you want to.

OpenStudy (waheguru):

but it is no revelent in this question right???

OpenStudy (jamesj):

It's not required, no.

OpenStudy (waheguru):

- im just in grade 9 new to this stuff

OpenStudy (waheguru):

thx fir ur help i was really confused

OpenStudy (jamesj):

...other than sense checking the answer. If either of the energy amounts we had found were less than 168kJ, then we would have known we had made an error.

OpenStudy (waheguru):

can u explain more plz

OpenStudy (waheguru):

so it had to be less than 168 because the input was 168??

OpenStudy (jamesj):

For the record, the energy efficiency of the microwave is Pragmatic work / Total work = 168 kJ / 336 kJ = 0.5 or 50% ====== I'm saying the total work done by the kettle or the microwave had to be at the very least equal to 168 kJ.

OpenStudy (jamesj):

...because if it had been less, then there is no way the 168 kJ of pragmatic work would have been realized.

OpenStudy (waheguru):

dont we divide output energy by input to find efficany

OpenStudy (jamesj):

In other words, nothing has more than 100% efficiency.

OpenStudy (jamesj):

What I'm calling "Pragmatic Work" is the output.

OpenStudy (jamesj):

The output is 168 kJ.

OpenStudy (waheguru):

oh so the output is 168 i was assuming that was the energy required being input

OpenStudy (jamesj):

Hence for the microwave, output = 168 kJ, input = 336 kJ. Therefore Efficiency = output/input = 168/336 = 0.5, or 50%

OpenStudy (waheguru):

ok thank-you so much i apperichate ur help if u want to see something cool and have google chrome tyoe tilt